### v1

# On the $\ell$-th largest degree of an intersecting family

Abstract
Let $\mathcal{F}\subset\binom{[n]}{k}$ be an intersecting family. For an element $i\in[n]$, the degree of $i$ is the number of sets in $\mathcal{F}$ that contain $i$. Assume that the degrees are ordered as $d_{1}\geq d_{2}\geq\cdots\geq d_{n}$. Huang and Zhao showed that if $n>2k$, then the minimum degree satisfies $d_{n}\leq\binom{n-2}{k-2}$, with the maximum attained by the $1$-star. We strengthen this result by proving that for $n\geq 2k+1$, the $(2k+1)$-th largest degree satisfies $d_{2k+1}\leq\binom{n-2}{k-2}$, thereby confirming a conjecture of Frankl and Wang. Furthermore, we prove that for $k>50$ and $n>\frac{11}{2}k$, the $(k+2)$-th largest degree $d_{k+2}$ is already at most $\binom{n-2}{k-2}$. The techniques we developed also yield an tight upper bound for the $(\ell+1)$-th largest degree $d_{\ell+1}$ for $\ell\leq k$ and sufficiently large $n$.
Keywords: Extremal set theory, Intersecting set family, Erdős–Ko–Rado theorem
1 Introduction
We denote by $[n]$ the standard $n$-element set $\{1,2,\dots,n\}$, $\binom{[n]}{k}$ the family of all $k$-element subsets of $[n]$, and $2^{[n]}$ by the power set of $[n]$. For a family $\mathcal{F}\subset 2^{[n]}$, we say $\mathcal{F}$ is $t$-intersecting if for any two sets $A,B\in\mathcal{F}$, we have $|A\cap B|\geq t$. If $\mathcal{F}$ is $1$-intersecting, we simply call it an intersecting family. If a $t$-intersecting family $\mathcal{F}$ is contained in $\binom{[n]}{k}$, we call $\mathcal{F}$ a $k$-uniform $t$-intersecting family. The famous Erdős–Ko–Rado theorem [3] states that if $\mathcal{F}$ is a $k$-uniform intersecting family, then for $n\geq 2k$, $|\mathcal{F}|\leq\binom{n-1}{k-1}$. Moreover, equality holds when $\mathcal{F}$ consists of all $k$-element sets containing a fixed element $i$, i.e. when $\mathcal{F}$ is isomorphic to a $1$-star. For this family, the element $i$ appears in $\binom{n-1}{k-1}$ members of $\mathcal{F}$, while every other element appears for $\binom{n-2}{k-2}$ times.
For a given family $\mathcal{F}\subset\binom{[n]}{k}$ and $i\in[n]$, denote by $d_{i}(\mathcal{F})$ the degree of $i$, defined as the number of sets in $\mathcal{F}$ that contain $i$, in other word, $d_{i}(\mathcal{F}):=|\{A|\,i\in A\in\mathcal{F}\}|.$ When the family $\mathcal{F}$ is clear from context and no confusion arises, we simply write $d_{i}(\mathcal{F})$ as $d_{i}$. Without loss of generality, we may assume that the degrees are ordered as $d_{1}\geq d_{2}\geq d_{3}\geq\dots\geq d_{n}$. In 2017, Huang and Zhao [7] used spectral graph theory to prove that in a $k$-uniform intersecting set family $\mathcal{F}$, the minimal degree $d_{n}(\mathcal{F})$ is at most $\binom{n-2}{k-2}$.
Theorem 1.1 (Huang, Zhao [7]).
Suppose $n>2k$ and $\mathcal{F}\subset\binom{[n]}{k}$ is an intersecting family. Then $d_{n}(\mathcal{F})\leq\binom{n-2}{k-2}$.
Recently, Frankl and Wang [4] observed that in a $k$-uniform intersecting family, the upper bound $\binom{n-2}{k-2}$ in fact governs not only the minimum degree but also most of the degrees. They proved that for n>6k9n>6k-9, the inequality $d_{2k+1}(\mathcal{F})\leq\binom{n-2}{k-2}$ holds. Moreover, they conjectured that the same bound should remain valid for all $n>2k$. In other words, every $k$-uniform intersecting family $\mathcal{F}\subset\binom{[n]}{k}$ satisfies $d_{2k+1}(\mathcal{F})\leq\binom{n-2}{k-2}$. If true, this conjecture would substantially strengthen the result of Huang and Zhao (Theorem 1.1).
In this paper, we verify their conjecture.
Theorem 1.2.
Let $n\geq 2k+1$ and let $\mathcal{F}\subset\binom{[n]}{k}$ be an intersecting family. Then
$d_{2k+1}(\mathcal{F})\leq\binom{n-2}{k-2}$.
A natural question is whether we can find more degrees that are bounded by $\binom{n-2}{k-2}$. The following extremal example of the Hilton–Milner theorem [6] provides a construction demonstrating that $d_{k+1}(\mathcal{F})\geq\binom{n-2}{k-2}+1$ may occur even for sufficiently large $n$.
Example 1.3 (Hilton–Milner).
Let $\mathcal{F}\subset\binom{[n]}{k}$ be the $k$-uniform intersecting set family consisting of the following subsets:
  1. (a)
    $\{2,3,\dots,k+1\}$;
  2. (b)
    all $k$-element sets that contain $1$ and intersect the set $\{2,3,\dots,k+1\}$.
Frankl and Wang [4] showed that for a $k$-uniform $t$-intersecting family $\mathcal{F}$, if $n>\binom{t+2}{2}k^{2}$, then $d_{k+2}(\mathcal{F})\leq\binom{n-t-1}{k-t-1}$. They also conjectured that the quadratic condition $n>\binom{t+2}{2}k^{2}$ can be strengthened to a linear range: there exists an absolute constant $c$ such that whenever $n>ckt$, any $k$-uniform $t$-intersecting family $\mathcal{F}$ satisfies $d_{k+2}(\mathcal{F})\leq\binom{n-t-1}{k-t-1}$. We verify their conjecture for $t=1$.
Theorem 1.4.
Let $k>50$, $n>\frac{11}{2}k$, and let $\mathcal{F}\subset\binom{[n]}{k}$ be an intersecting family. Then
$d_{k+2}(\mathcal{F})\leq\binom{n-2}{k-2}$.
Next we consider the $\ell$-th largest degree of an intersection family, Frankl and Wang [4] have determined the maximum possible values for $d_{\ell}(\mathcal{F})$ for $2\leq\ell\leq 4$. For the $(\ell+1)$-degree problem when $\ell\geq 4$, inspired by the Hilton–Milner construction, they provided the following example.
Example 1.5.
Let $\mathcal{F}$ be the $k$-uniform intersecting family consisting of all the following sets:
  1. (a)
    all $k$-element sets containing the element $1$ and at least one element from $\{2,3,\dots,\ell+1\}$;
  2. (b)
    all $k$-element sets containing the set $\{2,3,\dots,\ell+1\}$.
Then $d_{\ell+1}(\mathcal{F})=\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$.
They also prove that for $4\leq\ell\leq k$, $d_{\ell+1}\leq\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$ when $n>2\ell^{2}k$. In this paper, we obtain an improved range for large $\ell$ via the following theorem.
Theorem 1.6.
Suppose $20\leq\ell\leq k$, and $n>\frac{k^{2}}{\ell}+64k$. Then for every intersecting family $\mathcal{F}\subset\binom{[n]}{k}$, we have
$d_{\ell+1}(\mathcal{F})\leq\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$.
The remainder of the paper is organized as follows. In Section 2, we review the shifting methods that we will apply, and present several related lemmas. Section 3 contains the proof of Theorem 1.2. In Section 4, we employ similar techniques to establish both Theorem 1.4 and Theorem 1.6. Finally, Section 5 concludes with a summary of our results and two open questions.
2 The shifting method
In this section we give a brief description of the shifting method. This technique is particularly effective when dealing with intersecting families and will be employed in the proofs of all of our three theorems.
Definition 2.1.
For a given family $\mathcal{F}$ of subsets and $i,j\in[n]$ with $i\neq j$, the $ij$-shifting of a set $A\in\mathcal{F}$ is defined as:
$A_{ij}=\begin{cases}(A\cup\{j\})\setminus\{i\},&\text{if }i\in A,\ j\notin A,\ (A\cup\{j\})\setminus\{i\}\notin\mathcal{F};\\ A,&\text{otherwise}.\end{cases}$
The $ij$-shifting of the family $\mathcal{F}$ is defined by $\mathcal{F}_{ij}:=\{A_{ij}\mid A\in\mathcal{F}\}$.
Lemma 2.2 (Shifting lemma for intersecting family, [2]).
If $\mathcal{F}$ is a $t$-intersecting family, then $\mathcal{F}_{ij}$ is also a $t$-intersecting family.
The shifting lemma also holds for cross-intersecting families. We say two families $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are $t$-cross-intersecting if $|A\cap B|\geq t$ for all $A\in\mathcal{F}_{1}$ and $B\in\mathcal{F}_{2}$. We then have the following lemma:
Lemma 2.3 (Shifting lemma for cross-intersecting families, [2]).
If two families $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are $t$-cross-intersecting, then $(\mathcal{F}_{1})_{ij}$ and $(\mathcal{F}_{2})_{ij}$ are also $t$-cross-intersecting.
Definition 2.4 (Lexicographic order).
The lexicographic order is a total order $_{L}$ on $\binom{[n]}{k}$ defined by:
$A<_{L}B\quad\text{if and only if}\quad\min(A\setminus B)<\min(B\setminus A).$
The co-lexicographic order is a total order $_{C}$ on $\binom{[n]}{k}$ defined by:
$A<_{C}B\quad\text{if and only if}\quad\max(A\setminus B)<\max(B\setminus A).$
Theorem 2.5 ([1, 5]).
Suppose $\mathcal{F}_{1}\subset\binom{[n]}{a}$ and $\mathcal{F}_{2}\subset\binom{[n]}{b}$ are $t$-cross-intersecting families with $|\mathcal{F}_{1}|=x_{1}$ and $|\mathcal{F}_{2}|=x_{2}$. Let $\mathcal{L}(a,x_{1})$ denote the family consisting of the first $x_{1}$ $a$-sets in the lexicographic order, and similarly define $\mathcal{L}(b,x_{2})$. Then $\mathcal{L}(a,x_{1})$ and $\mathcal{L}(b,x_{2})$ are also $t$-cross-intersecting.
We note that the first $\binom{n-t}{a-t}$ $a$-sets in the lexicographic order on $\binom{[n]}{a}$ are exactly those containing the set $[t]$. Similarly, the first $\binom{n-t}{b-t}$ $b$-sets in the lexicographic order on $\binom{[n]}{b}$ are exactly those containing $[t]$. Moreover, when n>a+btn>a+b-t, for every set $A\in\binom{[n]}{a}$ that does not contain the whole set $[t]$, there exists a set $B$ with $t]\subset B\in\binom{[n]}{b}$ such that $|A\cap B|\leq t-1$. This observation yields the following corollary.
Corollary 2.6.
Under the same assumptions as in Theorem 2.5, if n>a+btn>a+b-t, then either $x_{1}\leq\binom{n-t}{a-t}$ or $x_{2}\leq\binom{n-t}{b-t}$.
Definition 2.7 (Shadow).
Let $\mathcal{F}\subset\binom{[n]}{k}$ be a $k$-uniform family. The shadow of $\mathcal{F}$, denoted by $\partial\mathcal{F}$, is the (k1)(k-1)-uniform family defined by:
$\partial\mathcal{F}:=\{A\mid A=B\setminus\{x\}\text{ for some }x\in B\in\mathcal{F}\}.$
Theorem 2.8 (Kruskal–Katona, [1, 8, 9]).
Let $\mathcal{F}\subset\binom{[n]}{k}$ with $|\mathcal{F}|=m$, and let $\mathcal{C}(k,m)$ denote the family of the first $m$ $k$-sets in the co-lexicographic order. Then the following inequality holds:
$|\partial\mathcal{F}|\geq|\partial\mathcal{C}(k,m)|.$
Corollary 2.9.
Let $\overline{\partial}\mathcal{F}$ denote the family
$\overline{\partial}\mathcal{F}=\{A\mid A=B\cup\{x\}\text{ for some }B\in\mathcal{F},\ x\in[n]\setminus B\}.$
Suppose $\mathcal{F}\subset\binom{[n]}{k}$ and $|\mathcal{F}|>\binom{n-1}{k-1}$. Then $|\overline{\partial}\mathcal{F}|>\binom{n-1}{k}$.
Proof.
Define $\mathcal{F}^{C}:=\{[n]\setminus A\mid A\in\mathcal{F}\}$. Then $|\mathcal{F}^{C}|=|\mathcal{F}|$, and
$\overline{\partial}\mathcal{F}=(\partial\mathcal{F}^{C})^{C}.$
Hence it suffices to prove that if $|\mathcal{F}^{C}|>\binom{n-1}{k-1}$, then $|\partial\mathcal{F}^{C}|>\binom{n-1}{k}$.
By Theorem 2.8, we may assume that $\mathcal{F}^{C}$ contains the first $\binom{n-1}{k-1}+1$ (nk)(n-k)-sets in the co-lexicographic order. These are all the (nk)(n-k)-sets contained in [n1][n-1], together with one additional set containing the element $n$. Consequently, $\partial\mathcal{F}^{C}$ must contain all (nk1)(n-k-1)-sets contained in [n1][n-1], plus at least one set containing $n$. Thus,
$|\partial\mathcal{F}^{C}|\geq\binom{n-1}{n-k-1}+1=\binom{n-1}{k}+1.$
Definition 2.10 ($\ell$-shifted family).
A $k$-uniform family $\mathcal{F}\subset\binom{[n]}{k}$ is called $\ell$-shifted if $\mathcal{F}_{ij}=\mathcal{F}$ for every pair $(i,j)$ with $i\notin[\ell]$ and $j\in[\ell]$.
Proposition 2.11.
Let $\mathcal{F}\subset\binom{[n]}{k}$ be a $t$-intersecting family and the degrees are ordered as $d_{1}(\mathcal{F})\geq d_{2}(\mathcal{F})\dots\geq d_{n}(\mathcal{F})$. Then there exists an $\ell$-shifted $t$-intersecting $k$-uniform family $\mathcal{F}^{\prime}$ such that $d_{i}(\mathcal{F}^{\prime})\geq d_{i}(\mathcal{F})$ for $i\leq\ell$. Furthermore, elements $\ell$] are still the $\ell$ elements with the largest degrees in $\mathcal{F^{\prime}}$. Consequently, the $\ell$-largest degree of $\mathcal{F}^{\prime}$ is not smaller than $d_{\ell}(\mathcal{F})$.
Proof.
If $\mathcal{F}$ is not $\ell$-shifted, then there exist indices $i\notin[\ell]$ and $j\in[\ell]$ such that $\mathcal{F}_{ij}\neq\mathcal{F}$. Replace $\mathcal{F}$ with $\mathcal{F}_{ij}$. By Lemma 2.2, $\mathcal{F}_{ij}$ remains $t$-intersecting. Note that the $ij$-shifting will not decrease the degree of $k\not=i$, so it won’t decrease the degree of $k\in[\ell]$. Therefore elements of $\ell$] give the $\ell$ largest degrees of $\mathcal{F^{\prime}}$, $d_{i}(\mathcal{F}_{ij})\geq d_{i}(\mathcal{F})\geq d_{\ell}(\mathcal{F})$ for $i\leq\ell$. We may repeatedly apply the $ij$-shifting for different pairs $(i,j)$ until this process terminates after finitely many steps, yielding the desired family $\mathcal{F}^{\prime}$. ∎
In the proof of our theorems, Proposition 2.11 allows us to assume that $\mathcal{F}$ is $\ell$-shifted. Furthermore, we can re-name the elements in the ground set so that the degrees are still in non-increasing order. The resulting family might not be $\ell$-shifted, but then we can apply shiftings again. Since at least one of the $\ell$ largest degrees strictly increases, while there are only finitely many subsets in the family, this process will eventually terminate, and we end up with a $\ell$-shifted family $\mathcal{F}$ whose degrees are in non-increasing order.
Proposition 2.12.
Let $\mathcal{F}$ be an $\ell$-shifted $t$-intersecting $k$-uniform family. For $A\in\mathcal{F}$, suppose there are elements $j\leq\ell$, $i>\ell$ such that $i\in A$ and $j\notin A$. Then $(A\cup\{j\})\setminus\{i\}\in\mathcal{F}$.
Proof.
Since $\mathcal{F}_{ij}=\mathcal{F}$ by the definition of $\ell$-shifted family, we have $A_{ij}\in\mathcal{F}_{ij}=\mathcal{F}$. Thus $(A\cup\{j\})\setminus\{i\}=A_{ij}\in\mathcal{F}$. ∎
Applying Proposition 2.12, we obtain the following property.
Lemma 2.13.
Let $\mathcal{F}$ be an $\ell$-shifted intersecting $k$-uniform family. Define the (not necessarily uniform) family $\mathcal{G}$ on $\ell$] by:
$\mathcal{G}=\{A\cap[\ell]\mid A\in\mathcal{F}\}.$
Then the following statements hold:
  1. (a)
    $\mathcal{G}$ is upward closed among subsets of $\ell$] of size at most $k$, meaning that for all $A\subset B\subset[\ell]$, if $A\in\mathcal{G}$ and $|B|\leq k$, then $B\in\mathcal{G}$.
  2. (b)
    If $\ell\geq 2k$, then $\mathcal{G}$ is intersecting.
  3. (c)
    If $A,B\in\mathcal{F}$ satisfy $A\cap B\cap[\ell]=\emptyset$, then
    $|A\cap[\ell]|+|B\cap[\ell]|+|A\cap B|\geq\ell+1.$
Proof.
(a) Suppose $A\in\mathcal{G}$ and $|A|<k$. Then there exists $A_{0}\in\mathcal{F}$ such that $A_{0}\cap[\ell]=A$. For any $j\in[\ell]\setminus A$, choose an element $i\in A_{0}\setminus A$ (note that $i\notin[\ell]$ since $A_{0}\cap[\ell]=A$). By Proposition 2.12,
$A_{1}:=(A_{0}\cup\{j\})\setminus\{i\}\in\mathcal{F}.$
Then $A_{1}\cap[\ell]=(A\cup\{j\})\cap[\ell]=A\cup\{j\}$. Hence $A\cup\{j\}\in\mathcal{G}$ for arbitrary $j\in[\ell]\setminus A$, proving that $\mathcal{G}$ is upward closed for sets of size at most $k$.
(b) Suppose for contradiction that $A,B\in\mathcal{F}$ with $A\cap B\cap[\ell]=\emptyset$. Then we know that $B\cap[\ell]\subset[\ell]\setminus A$. Choose a set $C$ satisfy $B\cap[\ell]\subset C\subset[\ell]\setminus A$ with $|C|=k$. By repeatedly applying Proposition 2.12, we obtain that $C\in\mathcal{F}$, but $C\cap A=\emptyset$, a contradiction.
(c) Assume $A,B\in\mathcal{F}$ with $A\cap B\cap[\ell]=\emptyset$. Let $C=A\cap B$ and suppose, for contradiction, that
$|C|\leq\ell-|A\cap[\ell]|-|B\cap[\ell]|.$
Then $|[\ell]\setminus(A\cup B)|=\ell-|A\cap[\ell]|-|B\cap[\ell]|\geq|C|$. Choose a set $D\subseteq[\ell]\setminus(A\cup B)$ with $|D|=|C|$. By repeatedly applying Proposition 2.12 to replace each element of $C$ in $A$ with a distinct element of $D$, we obtain that
$A^{\prime}:=(A\cup D)\setminus C\in\mathcal{F}.$
However, $A^{\prime}\cap B=\emptyset$, contradicting the assumption that $\mathcal{F}$ is intersecting. ∎
3 The $(2k+1)$-th largest degree is at most $\binom{n-2}{k-2}$
In this section we prove Theorem 1.2, which states that for any intersecting family $\mathcal{F}\subset\binom{[n]}{k}$ with $n>2k$, its $(2k+1)$-th largest degree satisfies $d_{2k+1}(\mathcal{F})\leq\binom{n-2}{k-2}$.
Proposition 2.11 allows us to assume that $\mathcal{F}$ is $(2k+1)$-shifted. Define the (not necessarily uniform) family $\mathcal{G}$ on $[2k+1]$ by:
$\mathcal{G}:=\{A\cap[2k+1]\mid A\in\mathcal{F}\}.$
Then part (b) of Lemma 2.13 guarantees that $\mathcal{G}$ is intersecting. Our general strategy is to find an element $i\in[2k+1]$ such that, for each $j\leq k$, $i$ appears in relatively few $j$-element sets of $\mathcal{G}$.
Definition 3.1 ($m$-degree).
Let $\mathcal{G}$ be a (not necessarily uniform) family of subsets of $[n]$. For $i\in[n]$, the $m$-degree of $i$ in $\mathcal{G}$, denoted by $d_{i}^{(m)}(\mathcal{G})$, is the number of sets in $\mathcal{G}$ of size $m$ that contain $i$:
$d_{i}^{(m)}(\mathcal{G}):=\bigl|\{A\in\mathcal{G}\mid i\in A,\ |A|=m\}\bigr|.$
We shall simply write $d_{i}^{(m)}$ when the family $\mathcal{G}$ is clear from the context.
Lemma 3.2.
Let $n>2k$ and let $\mathcal{G}$ be a (not necessarily uniform) intersecting family of subsets of $[n]$. Then there exists an element $i\in[n]$ such that for every $m\leq k$,
$d_{i}^{(m)}(\mathcal{G})\leq\binom{n-2}{m-2}$.
Proof.
Let $\overline{\mathcal{G}}$ denote the upward closure of $\mathcal{G}$:
$\overline{\mathcal{G}}=\{A\subseteq[n]\mid A\supseteq B\text{ for some }B\in\mathcal{G}\}.$
Then $\overline{\mathcal{G}}$ is still an intersecting family. Consequently, the $k$-element sets in $\overline{\mathcal{G}}$ form a $k$-uniform intersecting family on $[n]$ with $n>2k$. By Theorem 1.1, there exists an element $i\in[n]$ such that $d_{i}^{(k)}(\overline{\mathcal{G}})\leq\binom{n-2}{k-2}$.
We now prove by reverse induction on $m$ that
$d_{i}^{(m)}(\overline{\mathcal{G}})\leq\binom{n-2}{m-2}\quad\text{for all }m\leq k.$
The base case $m=k$ has already been established. Assume inductively that the statement holds for $m+1$ (where $m+1\leq k$), and suppose for contradiction that
$d_{i}^{(m)}(\overline{\mathcal{G}})>\binom{n-2}{m-2}$.
Define
$\overline{\mathcal{G}}_{i}^{(m)}:=\{A\setminus\{i\}\mid i\in A\in\overline{\mathcal{G}},\ |A|=m\}.$
This is an (m1)(m-1)-uniform family on the ground set $n]\setminus\{i\}$. Because $\overline{\mathcal{G}}$ is upward closed, we can apply Corollary 2.9 to $\overline{\mathcal{G}}_{i}^{(m)}$ on the ground set $n]\setminus\{i\}$ and obtain $|\,\overline{\partial}(\overline{\mathcal{G}}_{i}^{(m)})\,|>\binom{n-2}{m-1}$. Moreover, each $m$-set in $\overline{\partial}(\overline{\mathcal{G}}_{i}^{(m)})$ corresponds to an $(m+1)$-set in $\overline{\mathcal{G}}$ obtained by adding the element $i$, which yields
$d_{i}^{(m+1)}(\overline{\mathcal{G}})\geq|\,\overline{\partial}(\overline{\mathcal{G}}_{i}^{(m)})\,|>\binom{n-2}{m-1}$,
contradicting the induction hypothesis for $m+1$. Hence,
$d_{i}^{(m)}(\overline{\mathcal{G}})\leq\binom{n-2}{m-2}\quad\text{for all }m\leq k.$
Finally, since every set in $\mathcal{G}$ is contained in $\overline{\mathcal{G}}$, we have $d_{i}^{(m)}(\mathcal{G})\leq d_{i}^{(m)}(\overline{\mathcal{G}})$, which completes the proof. ∎
Proof of Theorem 1.2. By Proposition 2.11, we may assume that $\mathcal{F}$ is $(2k+1)$-shifted, and $d_{1},d_{2},\dots,d_{2k+1}$ are the $(2k+1)$ largest degrees.
Define $\mathcal{G}=\{A\cap[2k+1]\mid A\in\mathcal{F}\}$. By part (b) of Lemma 2.13, $\mathcal{G}$ is an intersecting family on $[2k+1]$. Applying Lemma 3.2 with $n=2k+1$, we obtain an element $i\in[2k+1]$ such that for every $m\leq k$,
$d_{i}^{(m)}(\mathcal{G})\leq\binom{2k-1}{m-2}$.
Observe that for a fix set $A\in\mathcal{G}$. There are at most $\binom{n-2k-1}{k-|A|}$ sets $B\in\mathcal{F}$ satisfying $B\cap[2k+1]=A$. Consequently, we can bound the degree of $i$ in $\mathcal{F}$ as follows:
$d_{i}(\mathcal{F})$ $\leq\sum_{m=1}^{k}d_{i}^{(m)}(\mathcal{G})\binom{n-2k-1}{k-m}$
$\leq\sum_{m=1}^{k}\binom{2k-1}{m-2}\binom{n-2k-1}{k-m}$
$=\binom{n-2}{k-2}$.
Hence, $d_{2k+1}(\mathcal{F})\leq d_{i}(\mathcal{F})\leq\binom{n-2}{k-2}$, which completes the proof. ∎
4 Estimating the $(\ell+1)$-th largest degree
In this section, we first prove Theorem 1.4, which states that for $k$ sufficiently large and $n>5.5k$, any intersecting family $\mathcal{F}\subset\binom{[n]}{k}$ satisfies $d_{k+2}(\mathcal{F})\leq\binom{n-2}{k-2}$. As we mentioned earlier, one cannot hope for $d_{k+1}(\mathcal{F})\leq\binom{n-2}{k-2}$ due to the Hilton–Milner example. Therefore to have the $1$-star as the extremal example, the index $k+2$ is best possible. We then apply similar techniques to prove Theorem 1.6, which determines the maximum value of $d_{\ell+1}(\mathcal{F})$ when $k\geq\ell\geq 20$ and $n$ is sufficiently large.
For a fixed $\ell$ and a subset $S\subseteq[\ell]$, we define the degree of $S$ (with respect to $\mathcal{F}$) as the size of the family
$d_{S}(\mathcal{F})=\bigl|\{A\in\mathcal{F}\mid A\cap[\ell]=S\}\bigr|.$
The following lemma shows that if two disjoint subsets $S_{1},S_{2}\subseteq[\ell]$ both have large degrees, a contradiction can be derived.
Lemma 4.1.
Let $\mathcal{F}$ be an $\ell$-shifted $k$-uniform intersecting family, and let $S_{1},S_{2}\subseteq[\ell]$ be disjoint subsets. Then either
$d_{S_{1}}(\mathcal{F})\leq\binom{n-\ell}{k-\ell-1+|S_{2}|}$
or
$d_{S_{2}}(\mathcal{F})\leq\binom{n-\ell}{k-\ell-1+|S_{1}|}$.
Proof.
For $X\subseteq[\ell]$, define
$\mathcal{F}(X):=\{A\setminus X\mid A\in\mathcal{F},\ A\cap[\ell]=X\}.$
For $i=1,2$, the family $\mathcal{F}(S_{i})$ is a $(k-|S_{i}|)$-uniform family on the ground set $n]\setminus[\ell]$. Take arbitrary $A\in\mathcal{F}(S_{1})$ and $B\in\mathcal{F}(S_{2})$. By part (c) of Lemma 2.13, since $S_{1}$ and $S_{2}$ are disjoint, we have,
$|A\cap B|\geq\ell+1-|S_{1}|-|S_{2}|.$
Hence, $\mathcal{F}(S_{1})$ and $\mathcal{F}(S_{2})$ are $(\ell+1-|S_{1}|-|S_{2}|)$-cross-intersecting.
Set $t:=\ell+1-|S_{1}|-|S_{2}|$. Applying Corollary 2.6 with parameters $a=k-|S_{1}|$, $b=k-|S_{2}|$, $t=t$, and ground set size $n$-$\ell$, we obtain that either
$|\mathcal{F}(S_{1})|\leq\binom{(n-\ell)-t}{a-t}\quad\text{or}\quad|\mathcal{F}(S_{2})|\leq\binom{(n-\ell)-t}{b-t}$.
Since the binomial coefficient $\binom{x}{y}$ is increasing in $x$ for $x\geq y\geq 0$, and $(n-\ell)-t\leq n$-$\ell$, we can relax both upper bounds to
$|\mathcal{F}(S_{1})|\leq\binom{n-\ell}{a-t}=\binom{n-\ell}{k-\ell-1+|S_{2}|}$,
$|\mathcal{F}(S_{2})|\leq\binom{n-\ell}{b-t}=\binom{n-\ell}{k-\ell-1+|S_{1}|}$.
respectively. Since $d_{X}(\mathcal{F})=|\mathcal{F}(X)|$ for any $X\subseteq[\ell]$, the conclusion follows. ∎
Proof of Theorem 1.4.
By Proposition 2.11, we may assume that $\mathcal{F}$ is $(k+2)$-shifted and that $d_{1}\geq d_{2}\geq\dots\geq d_{k+2}$ are the $k+2$ largest degrees. We are given the condition $n>\frac{11}{2}k$.
Set $m=\bigl\lfloor\frac{k}{2}\bigr\rfloor$. Define
$\mathcal{G}=\bigl\{A\subset[k+2]\mid|A|\leq m,\ d_{A}(\mathcal{F})>\tbinom{n-k-2}{m-3}\bigr\}$.
We claim that $\mathcal{G}$ is an intersecting family. Indeed, if there are two disjoint sets $A,B\in\mathcal{G}$, then Lemma 4.1 with $\ell=k+2$, $S_{1}=A$, $S_{2}=B$ yields either
$d_{A}(\mathcal{F})\leq\tbinom{n-k-2}{|B|-3}\quad\text{or}\quad d_{B}(\mathcal{F})\leq\tbinom{n-k-2}{|A|-3}$.
Since $|A|,|B|\leq m$ by definition of $\mathcal{G}$, we have $|A|-3,|B|-3\leq m-3$. Consequently, $\tbinom{n-k-2}{|B|-3}\leq\tbinom{n-k-2}{m-3}$ and $\tbinom{n-k-2}{|A|-3}\leq\tbinom{n-k-2}{m-3}$, because the binomial coefficient $\tbinom{N}{x}$ is increasing in $x$ for $0\leq x\leq N/2$ and m3m-3 is much smaller than (nk2)/2(n-k-2)/2 under the given condition $n>\frac{11}{2}k$. This contradicts the membership of $A$ and $B$ in $\mathcal{G}$. Hence $\mathcal{G}$ is intersecting.
Let $\overline{\mathcal{G}}:=\{A\subset[k+2]\mid A\supseteq B\text{ for some }B\in\mathcal{G}\}$ be the upward closure of $\mathcal{G}$. Then $\overline{\mathcal{G}}$ is also an intersecting family. Applying Lemma 3.2 with $n=k+2$ and $k$ replaced by $m$ (note that $2m<k+2$), we obtain an element $i\in[k+2]$ such that for every $u\leq m$,
$d_{i}^{(u)}(\overline{\mathcal{G}})\leq\binom{k}{u-2}$.
In particular, for $u=2$ we have $d_{i}^{(2)}(\overline{\mathcal{G}})\leq\binom{k}{0}=1$, therefore $d_{i}^{(2)}(\mathcal{G})\in\{0,1\}$. Below we discuss two cases according to the value of $d_{i}^{(2)}(\mathcal{G})$.
Case 1: $d_{i}^{(2)}(\overline{\mathcal{G}})=0$.
We split the set in $\mathcal{F}$ that contains $i$ into the following three parts and estimate the size of each part separately.
$D_{1}=\{A|i\in A\in\mathcal{F},|A\cap[k+2]|>m\},$
$D_{2}=\{A|i\in A\in\mathcal{F},|A\cap[k+2]|\leq m,A\cap[k+2]\not\in\mathcal{G}\},$
$D_{3}=\{A|i\in A\in\mathcal{F},A\cap[k+2]\in\mathcal{G}\}.$
(a) For $D_{1}$, consider $A\in\mathcal{F}$ with $i\in A$ and $|A\cap[k+2]|\geq m+1$. For each $j$ with $m+1\leq j\leq k$, there are at most $\binom{k+1}{j-1}$ $j$-element subsets $X\subseteq[k+2]$ containing $i$. For each such $X$, we have $d_{X}(\mathcal{F})\leq\binom{n-k-2}{k-j}$. Hence the size of $|D_{1}|$ is bounded by
$|D_{1}|\leq\sum_{j=m+1}^{k}\binom{k+1}{j-1}\binom{n-k-2}{k-j}\leq\Bigl(\sum_{j=m+1}^{k}\binom{k+1}{j-1}\Bigr)\binom{n-k-2}{k-m-1}\leq 2^{k+1}\binom{n-k-2}{k-m-1}$.
(b) For $D_{2}$, consider $A\in\mathcal{F}$ with $i\in A$, $|A\cap[k+2]|\leq m$, and $A\cap[k+2]\notin\mathcal{G}$. There are at most $2^{k+1}$ sets $X\subseteq[k+2]$ containing $i$. By definition of $\mathcal{G}$, each such $X$ satisfies $d_{X}(\mathcal{F})\leq\binom{n-k-2}{m-3}$. Thus,
$|D_{2}|\leq 2^{k+1}\binom{n-k-2}{m-3}$.
(c) For $D_{3}$, consider $A\in\mathcal{F}$ with $i\in A$ and $A\cap[k+2]\in\mathcal{G}$. Since $d_{i}^{(j)}(\mathcal{G})\leq d_{i}^{(j)}(\overline{\mathcal{G}})\leq\binom{k}{j-2}$ for $j\leq m$, and for each $j$-element set $X\in\mathcal{G}$ we have $d_{X}(\mathcal{F})\leq\binom{n-k-2}{k-j}$, we obtain (using $d_{i}^{(2)}(\overline{\mathcal{G}})=0$)
$|D_{3}|\leq\sum_{j=3}^{m}\binom{k}{j-2}\binom{n-k-2}{k-j}\leq\sum_{j=3}^{k}\binom{k}{j-2}\binom{n-k-2}{k-j}=\binom{n-2}{k-2}-\binom{n-k-2}{k-2}$.
Now we show that, under the conditions $n\geq 5.5k$ and $k\geq 50$,
$\binom{n-k-2}{k-2}\geq 2^{k+2}\binom{n-k-2}{k-m-1}>|D_{1}|+|D_{2}|.$
First, observe that
$\frac{\binom{n-k-2}{k-2}}{\binom{n-k-2}{k-m-1}}=\frac{\prod_{j=n-2k+1}^{\,n-2k+m-1}j}{\prod_{j=k-m}^{\,k-2}j}=\prod_{j=0}^{m-2}\frac{n-2k+j+1}{k-m+j}.$
When $n>5.5k$, each factor satisfies
$\frac{n-2k+j+1}{k-m+j}\geq\frac{n-2k+m-1}{k-2}>4.$
For $k\geq 50,n\geq 5.5k$, the following product is minimized when $n=5.5k$ and $k=50$. Direct computation then gives
$\prod_{j=0}^{9}\frac{n-2k+j+1}{k-m+j}\geq\left.\prod_{j=0}^{9}\frac{n-2k+j+1}{k-m+j}\right|_{k=50,n=5.5k}\geq 60\cdot 4^{10}.$
Consequently,
$\frac{\binom{n-k-2}{k-2}}{\binom{n-k-2}{k-m-1}}$ $=\prod_{j=0}^{m-2}\frac{n-2k+j+1}{k-m+j}$
$\geq\Biggl(\prod_{j=0}^{9}\frac{n-2k+j+1}{k-m+j}\Biggr)\Biggl(\prod_{j=10}^{m-2}\frac{n-2k+j+1}{k-m+j}\Biggr)$
[…]
$\geq 2^{k+2}.$
Finally, combining the three parts,
$d_{i}(\mathcal{F})=|D_{1}|+|D_{2}|+|D_{3}|<\binom{n-2}{k-2}$.
Since $i\in[k+2]$ and $d_{1},d_{2}\dots,d_{k+2}$ are the $k+2$ largest degrees, we have $d_{k+2}(\mathcal{F})\leq d_{i}(\mathcal{F})$. Hence,
$d_{k+2}(\mathcal{F})\leq\binom{n-2}{k-2}$.
Case 2: $d_{i}^{(2)}(\overline{\mathcal{G}})=1$.
Assume, without loss of generality, that $\{1,i\}\in\overline{\mathcal{G}}$. Because $\overline{\mathcal{G}}$ is upward closed, for any $j\leq m$ there are exactly $\binom{k}{j-2}$ $j$-element subsets of $[k+2]$ that contain $\{1,i\}$. Since $d_{i}^{(j)}(\overline{\mathcal{G}})\leq\binom{k}{j-2}$ for $j\leq m$ by Lemma 3.2, every $j$-element set in $\overline{\mathcal{G}}$ containing $i$ must also contain $1$.
Moreover, because $\overline{\mathcal{G}}$ is intersecting and $\{1,i\}\in\overline{\mathcal{G}}$, every set in $\overline{\mathcal{G}}$ must intersect $\{1,i\}$. Combined with the previous observation, it follows that every set in $\overline{\mathcal{G}}$ must contain the element $1$. Therefore, for every $j\neq 1$, we still have $d_{j}^{(u)}(\overline{\mathcal{G}})\leq\binom{k}{u-2}$ for all $u\leq m$ (the same bound as in the Lemma 3.2).
Now we discuss two subcases.
Subcase 2.1: There exists an element $j\neq 1$ such that $d_{j}^{(2)}(\overline{\mathcal{G}})=0$. In this situation, we can apply exactly the same argument as in Case 1 (with $j$ playing the role of $i$) to conclude that $d_{k+2}(\mathcal{F})\leq\binom{n-2}{k-2}$.
Subcase 2.2: For every $j\neq 1$, we have $d_{j}^{(2)}(\overline{\mathcal{G}})=1$. This means that $\{1,j\}\in\overline{\mathcal{G}}$ for all $j\in[k+2]\setminus\{1\}$. Then either $\{1\}\in\mathcal{G}$ or, for each $j\neq 1$, the set $\{1,j\}$ itself belongs to $\mathcal{G}$. We claim that every set in $\mathcal{F}$ contains the element $1$.
Now suppose, for contradiction, that there exists a set $A\in\mathcal{F}$ that does not contain the element $1$. We can pick an element $i\notin A\cap[k+2]$ with $i\neq 1$. Recall that $\mathcal{F}$ is $(k+2)$-shifted. By repeatedly applying Proposition 2.12, we can shift the set $A$ into $k+2]\setminus\{1,i\}$, and conclude that $A^{\prime}:=[k+2]\setminus\{1,i\}\in\mathcal{F}$. Because either $\{1,i\}\in\mathcal{G}$ or $\{1\}\in\mathcal{G}$, there exists a set $B\in\mathcal{F}$ such that $B\cap[k+2]=\{1,i\}\text{ or }\{1\}$. In either case we have $A^{\prime}\cap B=\emptyset$, a contradiction. Therefore, every set in $\mathcal{F}$ must contain the element $1$. Consequently,
$d_{k+2}(\mathcal{F})\leq\binom{n-2}{k-2}$.
This completes the proof for Case 2, and thus the proof of Theorem 1.4. ∎
Using similar techniques, we give the proof of Theorem 1.6.
Proof of Theorem 1.6:
In the following, we define the degree of a subset $S\subset[\ell+1]$ as
$d_{S}(\mathcal{F})=\bigl|\{A\in\mathcal{F}\mid A\cap[\ell+1]=S\}\bigr|.$
Suppose $n>\dfrac{k^{2}}{\ell}+64k$. By Lemma 2.11, we may assume that $\mathcal{F}$ is $(\ell+1)$-shifted.
Set $m=\bigl\lfloor\frac{\ell}{2}\bigr\rfloor$ and define
$\mathcal{G}=\bigl\{A\subset[\ell+1]\mid|A|\leq m,d_{A}(\mathcal{F})>\tbinom{n-\ell-1}{k-\ell-2+m}\bigr\}$.
Then, by Lemma 4.1 (applied with $\ell$ replaced by $\ell+1$), the (not necessarily uniform) family $\overline{\mathcal{G}}$ is intersecting. Applying Lemma 3.2 with $n=\ell+1$ and $k$ replaced by $m$, we obtain an element $i\in[\ell+1]$ such that $d_{i}^{(j)}(\overline{\mathcal{G}})\leq\binom{\ell-1}{j-2}$ for all $j\leq m$.
We consider the following two cases (if $d_{i}^{(2)}(\overline{\mathcal{G}})=1$, we assume without loss of generality that $\{1,i\}\in\overline{\mathcal{G}}$):
Case 1: $d_{i}^{(2)}(\mathcal{G})=0$ or $d_{\{1,i\}}(\mathcal{F})\leq\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$.
We split the set in $\mathcal{F}$ that contains $i$ into three parts and estimate the size of each part separately.
$D_{1}=\{A|i\in A\in\mathcal{F},|A\cap[\ell+1]|>m\},$
$D_{2}=\{A|i\in A\in\mathcal{F},|A\cap[\ell+1]|\leq m,A\cap[\ell+1]\not\in\mathcal{G}\},$
$D_{3}=\{A|i\in A\in\mathcal{F},A\cap[\ell+1]\in\mathcal{G}\}.$
(a) For $D_{1}$, consider those $A$ satisfy $i\in A\in\mathcal{F}$ and $|A\cap[\ell+1]|\geq m+1$. Similar to the proof of Theorem 1.4, we bound the number of $|D_{1}|$ by:
$|D_{1}|\leq\sum_{j=m+1}^{\ell+1}\binom{\ell}{j-1}\binom{n-\ell-1}{k-j}\leq 2^{\ell}\binom{n-\ell-1}{k-m-1}$.
(b) For $D_{2}$, those $A$ satisfy $i\in A\in\mathcal{F}$, $|A\cap[\ell+1]|\leq m$, but $A\cap[\ell+1]\notin\mathcal{G}$. The number of such $A$ can be bounded by:
$|D_{2}|\leq 2^{\ell}\binom{n-\ell-1}{k-\ell-2+m}$.
(c) For $D_{3}$, consider those $A$ with $i\in A\in\mathcal{F}$ with $A\cap[\ell+1]\in\mathcal{G}$. Then:
$|D_{3}|$ $\leq d_{\{1,i\}}(\mathcal{F})+\sum_{j=3}^{m}\binom{\ell-1}{j-2}\binom{n-\ell-1}{k-j}$
$\leq\binom{l-1}{0}\binom{n-l-1}{k-2}-\binom{n-k-2}{k-2}+\sum_{j=3}^{k}\binom{\ell-1}{j-2}\binom{n-\ell-1}{k-j}$
$=\binom{n-2}{k-2}-\binom{n-k-2}{k-2}$.
Now we show that, under the conditions $n>\dfrac{k^{2}}{\ell}+64k$ and $k\geq\ell\geq 20$,
$\binom{n-k-2}{k-2}\geq 2^{\ell+1}\binom{n-\ell-1}{k-m-1}>|D_{1}|+|D_{2}|.$
$\frac{\binom{n-k-2}{k-2}}{\binom{n-\ell-1}{k-m-1}}$ $=\prod_{j=1}^{k-m-1}\frac{n-k-1-j}{n-\ell-j}\;\prod_{j=1}^{m-1}\frac{n-2k+j}{k-m-1+j}$
$\geq\left(1-\frac{\ell}{k+62\ell}\right)^{k-m-1}\left(\frac{k}{\ell}+62\right)^{m-1}$
[…]
$\geq 2^{\ell+1}.$
The first inequality is because
$\dfrac{n-k-1-j}{n-\ell-j}=1-\dfrac{k+1-\ell}{n-\ell-j}\geq 1-\dfrac{k}{\frac{k^{2}}{\ell}+64k-\ell-k}\geq 1-\frac{\ell}{k+62\ell},$
$\dfrac{n-2k+j}{k-m-1+j}\geq\dfrac{\frac{k^{2}}{\ell}+64k-2k+m-1}{k-2}\geq\dfrac{k}{\ell}+62$
Therefore,
$d_{i}(\mathcal{F})\leq|D_{1}|+|D_{2}|+|D_{3}|\leq\binom{n-2}{k-2}<\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$.
Case 2: $d_{i}^{(2)}(\mathcal{G})=1$ and $d_{\{1,i\}}(\mathcal{F})>\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$. Then, by the same argument as in the proof of Theorem 1.4, we have $d_{j}^{(u)}(\overline{\mathcal{G}})\leq\binom{\ell-1}{u-2}$ for all $j\neq 1$ and $u\leq m$. Consequently, if for some $j\neq 1$ we have $d_{\{1,j\}}(\mathcal{F})\leq\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$, then the proof is already complete (by reducing to the situation of Case 1). Thus we may assume that for every $j\neq 1$,
$d_{\{1,j\}}(\mathcal{F})>\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$.
We shall show that under this assumption every set in $\mathcal{F}$ must either contain the element $1$ or contain the whole set $\{2,3,\dots,\ell+1\}$.
Assume, for contradiction, that there exists a set $A\in\mathcal{F}$ that not contains $1$ and $i$ with $2\leq i\leq\ell+1$. By part (a) of Lemma 2.13 (upward‑closedness), we can find a set $B\in\mathcal{F}$ such that $B\cap[\ell+1]=[\ell+1]\setminus\{1,i\}$. Then $|B\setminus[\ell+1]|=k-\ell+1$.
Now consider the family
$\mathcal{F}(\{1,i\}):=\{A\setminus\{1,i\}\mid A\cap[\ell+1]=\{1,i\},A\in\mathcal{F}\}.$
Every member of $\mathcal{F}(\{1,i\})$ must intersect $B\setminus[\ell+1]$ (otherwise the corresponding original set $A$ would be disjoint from $B$). Hence
$|\mathcal{F}(\{1,i\})|\leq\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$.
But $|\mathcal{F}(\{1,i\})|=d_{\{1,i\}}(\mathcal{F})$, which contradicts our assumption that $d_{\{1,i\}}(\mathcal{F})>\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$. Therefore, in this case every set in $\mathcal{F}$ either contains $1$ or contains $\{2,3,\dots,\ell+1\}$. Consequently,
$d_{\ell+1}(\mathcal{F})\leq\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$.
5 Concluding Remarks
In this paper, we study the problem of determining the maximum value of the $(\ell+1)$-th largest degree of an intersecting family $\mathcal{F}\subset\binom{[n]}{k}$. For $\ell=2k$, Theorem 1.2 provides the precise range that the $1$-star gives the maximum, resolving the Frankl–Wang Conjecture. For $\ell=k+1$, Theorem 1.4 shows that the $1$-star example gives the maximum for $n>\frac{11}{2}k$. A natural question is whether the constant $\frac{11}{2}$ can be reduced to $2$.
Question 5.1.
Can we find a constant $M$ such that for every $n\geq 2k+M$, every intersecting family $\mathcal{F}\subset\binom{[n]}{k}$ satisfies
$d_{k+2}(\mathcal{F})\leq\binom{n-2}{k-2}$?
In particular, is $M=3$ already sufficient?
As a remark to Question 5.1, we present the following examples for $n=2k+2$ and $k\in\{4,5,7\}$. In all these examples, there exists an intersecting family whose $(k+2)$-th degree is strictly greater than $\binom{n-2}{k-2}$.
Example 5.2.
Let $k=4$, $n=10$. Construct $\mathcal{F}$ as the union of the following two families:
  1. (a)
    all $4$-element subsets of $[6]$, i.e., $\binom{[6]}{4}$;
  2. (b)
    all sets of the form $A\cup\{b\}$, where $A$ is a block of a $2\text{-}(6,3,2)$ design and $b\in\{7,8,9,10\}$.
One can verify that $\mathcal{F}$ is intersecting and that the $6$th largest degree in $\mathcal{F}$ equals $30$, which exceeds $\binom{n-2}{k-2}=\binom{8}{2}=28$.
Example 5.3.
For $k=5$ and $n=12$, take a Fano plane on $[7]$, denoted by $\mathcal{P}$ (a Steiner triple system $S(2,3,7)$). Let $\mathcal{F}$ consist of all $5$-element subsets of $[12]$ that contain at least one triple from $\mathcal{P}$. It can be checked that $\mathcal{F}$ is intersecting under this construction, and that $d_{7}(\mathcal{F})=125>\binom{10}{3}=120$.
Example 5.4.
For $k=7$ and $n=16$, we let the ground set be $[16]$ and let $\mathcal{F}$ consists of all the $7$-sets whose intersection with $[9]$ contains a subset from the family $\mathcal{G}=\{S\cup\{i\}:S\in\mathcal{H},i\in\{7,8,9\}\}\cup\{\{j,7,8,9\}:j\in[6]\}$. Here $\mathcal{H}$ is the unique 22-$(6,3,2)$ desing on $[6]$. Counting gives $d_{9}(\mathcal{F})=2023>2002=\binom{n-2}{k-2}$.
Additionally, when $n=2k+1$ for an odd integer $k\geq 3$, we may take $\mathcal{F}$ to consist of all $k$-sets that intersect $[k+2]$ in at least $\frac{k+3}{2}$ elements. It is straightforward to check that $d_{k+2}(\mathcal{F})>\binom{n-2}{k-2}$. We omit the calculations here.
Regarding the $(\ell+1)$-th largest degree with $\ell\leq k$, our Theorem 1.6 shows that when $\ell=\Theta(k)$, the bound $d_{\ell+1}\leq\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$ holds for $n>\Omega(k)$. Since Frankl and Wang showed that the same bound is valid for $\ell=\Omega(1)$, a natural question is whether a unified statement can be made for all $\ell$.
Question 5.5.
Can we find an absolute constant $C$ such that for every $4\leq\ell\leq k$ and every $n>Ck$, any intersecting family $\mathcal{F}\subset\binom{[n]}{k}$ (with degrees ordered as $d_{1}\geq d_{2}\geq\cdots\geq d_{n}$) satisfies
$d_{\ell+1}(\mathcal{F})\leq\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$?
If such a constant exists, what is the smallest possible value of $C$?